OpenStudy (anonymous):

what is the value of the expression-x(y-8)^2

jimthompson5910 (jim_thompson5910):

\[x(y-8)^2\] \[x(y-8)(y-8)\] \[x(y^2-16y+64)\] \[xy^2-16xy+64x\] So \[x(y-8)^2=xy^2-16xy+64x\]

OpenStudy (anonymous):

oops i forgot if the value of x is -2 and the value of y is 5

jimthompson5910 (jim_thompson5910):

\[x(y-8)^2\] \[(-2)(5-8)^2\] \[(-2)(-3)^2\] \[(-2)(9)\] \[-18\] So \[x(y-8)^2=-18\] when x=-2 and y=5

OpenStudy (anonymous):

thanks